# Refining Skewed Perceptions in Vision-Language Models through Visual Representations

### v1

@@ Abstract
Large vision-language models (VLMs), such as CLIP, have become foundational, demonstrating remarkable success across a variety of downstream tasks. Despite their advantages, these models, akin to other foundational systems, inherit biases from the disproportionate distribution of real-world data, leading to misconceptions about the actual environment. Prevalent datasets like ImageNet are often riddled with non-causal, spurious correlations that can diminish VLM performance in scenarios where these contextual elements are absent. This study presents an investigation into how a simple linear probe can effectively distill task-specific core features from CLIP’s embedding for downstream applications. Our analysis reveals that the CLIP text representations are often tainted by spurious correlations, inherited in the biased pre-training dataset. Empirical evidence suggests that relying on visual representations from CLIP, as opposed to text embedding, is more practical to refine the skewed perceptions in VLMs, emphasizing the superior utility of visual representations in overcoming embedded biases. Our codes will be available in here.
@@ 1 Introduction
Models like Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) have set new benchmarks across various tasks by contrastively matching semantically closest image and text pairs. However, due to the disproportionate distribution embedded in real-world datasets like ImageNet (Deng et al., 2009) or LAION (Schuhmann et al., 2021), pre-trained VLMs inherently acquire biases from these large-scale datasets. This phenomenon, known as spurious correlation, refers to patterns that correlate the target class with non-causal contextual elements. For instance, a vision model may classify cows correctly but fail when cows appear outside the typical grassland background, revealing grass as a shortcut predictor for cow (Beery et al., 2018). Similarly, BERT’s (Devlin et al., 2018) peak performance on the argument reasoning comprehension task is largely due to exploiting spurious statistical cues in the dataset, like the negation word “not” (Niven and Kao, 2019).
@@
To answer these questions, we conduct various experiments. First, we assess CLIP’s zero-shot learning performance on the widely used Waterbirds dataset (Sagawa et al., 2019) before and after removing the “background” context. Next, we explore the expressiveness limits of CLIP embeddings by performing various classification tasks on the CelebA (Liu et al., 2018) dataset using only linear probing to see if the embeddings capture nuanced features. To examine the practicality of zero-shot classification in the presence of spurious correlations, we evaluate the degree of contamination in CLIP’s text representations due to biased pre-training data through extensive statistical analysis. Futher, to investigate the language’s ability to guide the path to optimal linear probe, we develop a framework called PromptCraft, which recovers human-readable text from embeddings derived by image/text encoders or even linear layer parameters. Lastly, we show CLIP’s visual representation’s ability to distill core features using the proposed VisualDistiller framework.
@@
We develop a simple yet effective text recovery framework called PromptCraft that recovers text from vector embeddings. We find that CLIP’s text embeddings are contaminated by diverse elements, making text embeddings impractical for debiasing the model.
We demonstrate that using visual embeddings from CLIP to distill visual representations is highly effective. The debiased features achieve excellent performance in group accuracy comparable to supervised methods like DFR, which offers a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
@@ 2 Related Work
Deep learning frameworks frequently exhibit uneven performance across various groups due to spurious correlations, resulting in notably lower test accuracy for minority groups compared to majority groups. This issue contrasts with the training phase, where both groups generally achieve more balanced training accuracy (Sagawa et al., 2019; Geirhos et al., 2020). (Geirhos et al., 2020; Shah et al., 2020; Hermann and Lampinen, 2020) highlights that neural networks are prone to a simplicity bias, often emphasizing trivial spurious features while neglecting the essential core features.
To address these challenges, substantial research has been dedicated to enhancing robustness against spurious correlations. When group labels are available, strategies such as class balancing (He and Garcia, 2009; Cui et al., 2019), importance weighting (Shimodaira, 2000; Byrd and Lipton, 2019), robust optimization (Sagawa et al., 2019; Kirichenko et al., 2022; Izmailov et al., 2022), and contrastive learning (Taghanaki et al., 2021) have been developed to ensure balanced training across different group sizes. In scenarios where group labels are unavailable, a common approach involves initially training an auxiliary model using empirical risk minimization (ERM). The predictions from this model are then used to infer group information, which in turn guides the training of a more robust second model. This robust model is typically trained using techniques such as sample balancing (Liu et al., 2021; Nam et al., 2020), or contrastive learning (Zhang et al., 2022; Zhang and Ré, 2022; Yang et al., 2023) with the inferred group labels.
@@
VLMs have gained increased popularity for their ability to perceive the world through multiple modalities. Previous research has sought to enhance the robustness of vision classifiers by incorporating language features, using techniques such as attention maps (Petryk et al., 2022) and modifications to feature attributes (Zhang et al., 2023). Significant advancements (Yang et al., 2023; Zhang and Ré, 2022) have been made in developing pre-trained multimodal models resistant to spurious correlations. For instance, (Zhang and Ré, 2022) proposes a novel contrastive adapter that, when combined with transfer learning, improves group robustness. However, this method does not always lead to better results, especially for specific downstream applications. Conversely, (Yang et al., 2023) pioneers a fine-tuning strategy specifically designed to address spurious correlations with group labels in pre-trained multimodal models. (Chuang et al., 2023) addresses VLMs’ bias in zero-shot classification by projecting out biased directions in the text embeddings. Unlike these approaches, our objective is to investigate the inherent skewed perception embedded in all text embeddings (including target class text and spurious attribute text) and explore the possibilities of using visual representations instead to distill the task-specific core features from VLMs like CLIP for downstream tasks, without the need for group annotations.
@@ 3 Preliminaries
In this study, we explore the spurious features inherent in the CLIP visual representations and assess their impact on classification performance. For a given classification task, we have samples , where represents the input features, as the class labels, as the spurious attributes, and as the group labels. We examine scenarios of distribution shifts occurring between samples across different groups but within the same class. In the Waterbirds dataset (Sagawa et al., 2019), we define {landbird, waterbird}, {land background, water background}, and {landbird on land (), landbird on water (), waterbird on land (), waterbird on water ()}. Notably, and are the minority groups (fewer training samples), whereas and are the majority groups. In the default CelebA dataset (Liu et al., 2018), the categories are {non-blond hair, blond hair}, {female, male}, and {non-blond hair female (), blond hair female (), non-blond hair male (), blond hair male ()}. With regard to CelebA, are the majority groups, with being the minority group.
@@
4 Does there exist a task matrix that can achieve optimal task performance?
In modern machine learning system design, the goal is often to enhance foundational models with specialized modules for specific downstream tasks. For classification tasks in particular, the ideal is to utilize the same representations derived from a pre-trained model across various classification challenges. Given the nature of spurious correlations — where a feature deemed spurious for one task may be essential for another — we expect the VLMs to capture a broad spectrum of nuanced visual information, and removing the spurious feature by specialized modules. This section delves into the presence of spurious correlations within VLMs and explores whether a simple linear probe — referred to here as the task matrix — can deliver optimal performance.
4.1 Unraveling Spurious Correlations in Vision-Language Models
In this section, we first try to investigate the existence of spurious features within VLMs through a comparative analysis of zero-shot classification performance on the Waterbirds dataset. To show that spurious correlation impairs the performance of VLMs, we conduct two experiments: zero-shot classification using the original Waterbirds dataset (with natural background) and using a modified version of the Waterbirds dataset from which the background have been erased based on mask.
Figure 1 presents the group accuracy changes across these two scenarios. In the first scenario (using original data), the models exhibit uneven accuracies across majority and minority groups, reflecting the unbalanced group robustness across the dataset. When the backgrounds are removed, the accuracy in recognizing and (minority groups) boosts, as the confounding elements are no longer able to mislead the model. Despite the relatively small change on majority , we see consistent accuracy drop in majority , across three different architectures. The disparity in performance and drastic accuracy change in both minority and majority group, confirms our hypothesis that visual representations in current models are entangled with spurious features that significantly impair classification performance. This raises an foundational question: does this imply that we are unable to achieve flawless task execution on CLIP representations when faced with spurious features? If not, how?
@@ 4.2 Assessing the Expressiveness of CLIP’s Visual Representations under Linear Probing
In order to see the upper limit of CLIP visual representation with linear transformation, we applied deep feature reweighting (DFR) (Kirichenko et al., 2022) to 29 attribute classification challenges (see Table 3 for details) on the CelebA dataset, where each attribute demonstrated a gender-biased distribution. Likewise, they also involve four groups based on gender and attribute presence: female without [attribute] (), male without [attribute] (), female with [attribute] (), and male with [attribute] (). Following DFR’s implementation, a linear layer is attached to the CLIP image encoder to facilitate binary classification, with updates restricted solely to the weights of the linear layer. As a supervised method, DFR usually signifies the peak performance that a linear layer can attain, by strategically adjusting sample weights based on their group frequency to enhance accuracy, particularly for underperforming groups, and reduce the impact of spurious attributes without altering the primary network. High accuracy in these groups suggests that the standard CLIP image encoder successfully captures essential task-related features, not merely relying on these features for predictions under ERM.
Figure 2 showcases the outcomes for various attribute classifications on CLIP ViT-L/14 model, sorted by ascending average group accuracy. The spectrum of attributes ranged from subtle features like straight hair and narrow eyes to more overt characteristics such as eyeglasses, baldness, and hats. Notably, the DFR approach strategy enabled a majority of the attributes — over 25 out of 29 — to achieve more than 75% worst group accuracy (WGA), with more than 23 attribute classification tasks surpassing 80% average group accuracy. The five attributes with the highest accuracies exceeded 95% in both the worst and average group accuracy measures. These results underscore how fine-grained the CLIP’s visual representations are, capturing a comprehensive spectrum of visual information, including subtle features that are typically challenging for human perception. Hence, we believe that visual representations learned by CLIP are adept at extracting nuanced features within images for various tasks by linear transformation.
5 Can language unveil the path to the optimal task matrix?
VLMs like CLIP prevail partly because of their capability to perform zero-shot inferences guided by intuitive language cues. As demonstrated in the previous section, techniques such as DFR guide us toward identifying an optimal task matrix that can effectively discern variations in core features. Similarly, in zero-shot learning, the task matrix is formed by concatenating text representations. This section explores the biases inherent in zero-shot classification prompts and examines whether it is possible to find a text prompt whose representation vector closely approximates the optimal task matrix, akin to a linear probe refined by DFR.
5.1 Language Representations are not as Pristine as One Might Thought
@@
To test this hypothesis, we examined various prompts and corresponding image pairs, calculating the cosine similarity between them. For instance, we evaluated pairs like (“a photo of a camel”/“a photo of a cow”, desert/pasture images), (“a photo of a polar bear”/“a photo of a panda”, glacier/bamboo forest images), and conversely (“a photo of a pasture”/“a photo of a desert”, camel/cow images), (“a photo of bamboo forest”/“a photo of glacier”, polar bear/panda images), with ensuring the images tested here did not contain the objects mentioned in the prompts. This methodology helps quantify the extent of spurious features embedded in text representations.
@@
Figure 3 shows the result. Notably, the cosine similarity distributions, indicated by the mean (green triangle) and median (orange line), reveal strong correlations—for example, the prompt “a photo of a camel” with camel-free desert images and the prompt “a photo of a cow” with cow-free pasture images. This pattern is consistent across various tested pairs, underscoring the substantial presence of context-related features in CLIP text representations that are not explicitly present in the prompts. For the prompt and image pair with less pronounced correlations, like “a photo of a dog” to forest and desert, we did not observe the same level of disparity in mean and median value, see Figure 6 in Appendix for more details.
@@
These results suggest that using text representations for zero-shot classification or debiasing with the representations from spurious attribute prompts (Chuang et al., 2023) could lead to unexpected outcomes, due to the embedded non-target features.
5.2 Can we optimize text prompts through inverse problem solving?
The results illustrated in Figure 2 show that leveraging DFR can recover an optimal task matrix for binary classification. This raises the question: is it possible to identify a text prompt whose representation closely approximates the DFR-trained linear probe? If achievable, this may give us clue of how to reach equivalent performance through zero-shot classification using textual prompting in the future.
In this section, we introduce a text recovery workflow named PromptCraft, a framework designed to identify a text prompt whose representation aligns with a specified target vector, sourced from either CLIP’s image/text encoder or even the weights of a linear layer, as depicted in Figure 7. The entire framework focuses on optimizing the token embeddings, , which is the only learnable tensor in the framework. We initiate the process with a initial text such as “a photo of dog”, anticipating that the final recovered text will follow the format “a photo of [object]”, thereby simplifying our optimization approach by starting with a prompt close to the desired outcome. After passing the initialized by initial text through the frozen CLIP text encoder, we extract the end-of-text vector, , as our resultant text representation. The similarity between and the target vector is measured using the designated loss function
which guides the backward propagation to refine . Upon convergence of the loss, is mapped back to the tokens most similar in terms of cosine similarity. Finally, the recovered tokens are decoded into a human-readable text prompt using the CLIP token decoder. Reader can refer to Algorithm 1 and Figure 7 for better understanding.
To validate the effectiveness of PromptCraft, we performed 4,537 experiments and show the results in Figure 4. For each experiment, the target vectors are from using CLIP to encode the target text (e.g., “a photo of a cat”). We initialize the learnable embedding with the token embedding of various initial text (e.g., “a photo of a dog”), as random initialization may add extra difficulty to the recovery. Although the final might not exactly match , we considered a trial successful if the text decoded from the learned token embeddings contained the target text, and unsuccessful otherwise.
In Figure 4, every cell’s color in the matshow suggests whether the experiment on the pair (initial text, target text) succeed. The cell filled with green color means that the recovery is successfully performed on the text pair. In the left matshow, a total of 1,936 animal related text pairs were tested, with 82% of the experiment succeed. 100% of the animal text can be recovered from at least 26 initial text. In the right matshow, a total of 2,601 food related text pairs were tested, with 80% of the experiment succeed. 100% of the food text can be recovered from at least 21 initial text. This demonstrates the efficacy of our PromptCraft framework in reconstructing text from vectors, robust to different start point or end point in the feature space.
Since we posit that the linear probe, trained via DFR, precisely captures the core features pertinent to our task, we employed the DFR-trained linear probe as and utilized PromptCraft to approximate the ideal human-readable text prompt. We showcase the recovered text from DFR linear probe (landbird vector) in Table 1 with various EOT position. Despite these efforts, the results below suggest that isolating core from spurious features using an optimal human-readable text prompt still remains a challenging endeavor. For more result, please refer to Table 8 and Table 9 in Appendix.
| EOT Index | Recovered Text |
| 6 | tman photo hyuk landbird |
| 7 | ta mascot lowers !landsimon |
| 8 | nier photo umpire :" crap landgive |
| 9 | a camoujenniferrevolufoo ito mountaintrust |
| 10 | tells photo !! birds ent rattwouldn |
6 Can image unveil the path to the optimal task matrix?

### v2

@@ Abstract
Large vision-language models (VLMs), such as CLIP, have become foundational, demonstrating remarkable success across a variety of downstream tasks. Despite their advantages, these models, akin to other foundational systems, inherit biases from the disproportionate distribution of real-world data, leading to misconceptions about the actual environment. Prevalent datasets like ImageNet are often riddled with non-causal, spurious correlations that can diminish VLM performance in scenarios where these contextual elements are absent. This study presents an investigation into how a simple linear probe can effectively distill task-specific core features from CLIP’s embedding for downstream applications. Our analysis reveals that the CLIP text representations are often tainted by spurious correlations, inherited in the biased pre-training dataset. Empirical evidence suggests that relying on visual representations from CLIP, as opposed to text embedding, is more practical to refine the skewed perceptions in VLMs, emphasizing the superior utility of visual representations in overcoming embedded biases. Our codes will be available upon publication.
@@ 1 Introduction
VLMs like Contrastive Language-Image Pre-training (CLIP) [17] have set new benchmarks across various tasks by contrastively matching semantically closest image and text pairs. However, due to the disproportionate distribution embedded in real-world datasets like ImageNet [5] or LAION [19], pre-trained VLMs inherently acquire biases from these large-scale datasets. This phenomenon, known as spurious correlation, refers to patterns that correlate the target class with non-causal contextual elements. For instance, a vision model may classify cows correctly but fail when cows appear outside the typical grassland background, revealing grass as a shortcut predictor for cow [1]. Similarly, BERT’s [6] peak performance on the argument reasoning comprehension task is largely due to exploiting spurious statistical cues in the dataset, like the negation word “not” [15].
@@
To answer these questions, we conduct various experiments. First, we assess CLIP’s zero-shot learning performance on the widely used Waterbirds dataset [18] before and after removing the “background” context. Next, we explore the expressiveness limits of CLIP embeddings by performing various classification tasks on the CelebA [13] dataset using only linear probing to see if the embeddings capture nuanced features. To examine the practicality of zero-shot classification in the presence of spurious correlations, we evaluate the degree of contamination in CLIP’s text representations due to biased pre-training data through extensive statistical analysis. Lastly, we show CLIP’s visual representation’s ability to distill core features using the proposed VisualDistiller framework.
@@
We find that CLIP’s text embeddings are contaminated by diverse elements, making text embeddings impractical for debiasing the model.
We demonstrate that using visual embeddings from CLIP to distill visual representations is highly effective. The debiased features achieve excellent performance in group accuracy comparable to supervised methods like DFR [11], which offers a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
@@ 2 Related Work
Deep learning frameworks frequently exhibit uneven performance across various groups due to spurious correlations, resulting in notably lower test accuracy for minority groups compared to majority groups. This issue contrasts with the training phase, where both groups generally achieve more balanced training accuracy [18, 7]. [7, 20, 9] highlights that neural networks are prone to a simplicity bias, often emphasizing trivial spurious features while neglecting the essential core features.
To address these challenges, substantial research has been dedicated to enhancing robustness against spurious correlations. When group labels are available, strategies such as class balancing [8, 4], importance weighting [21, 2], robust optimization [18, 11, 10], and contrastive learning [22] have been developed to ensure balanced training across different group sizes. For example, deep feature reweighting (DFR) [11] solves this by retraining the last linear layer of the model (e.g. ResNet) by upweighting the underrepresented samples in the loss function, ensuring better handling of minority groups. In scenarios where group labels are unavailable, a common approach involves initially training an auxiliary model using empirical risk minimization (ERM). The predictions from this model are then used to infer group information, which in turn guides the training of a more robust second model. This robust model is typically trained using techniques such as sample balancing [12, 14], or contrastive learning [27, 26, 24] with the inferred group labels.
@@
VLMs have gained increased popularity for their ability to perceive the world through multiple modalities. Previous research has sought to enhance the robustness of vision classifiers by incorporating language features, using techniques such as attention maps [16] and modifications to feature attributes [28]. Significant advancements [24, 26] have been made in developing pre-trained multimodal models resistant to spurious correlations. For instance, [26] proposes a novel contrastive adapter that, when combined with transfer learning, improves group robustness. However, this method does not always lead to better results, especially for specific downstream applications. Conversely, [24] pioneers a fine-tuning strategy specifically designed to address spurious correlations with group labels in pre-trained multimodal models. [3] addresses VLMs’ bias in zero-shot classification by projecting out biased directions in the text embeddings. [25] proposed composition-aware hard negatives during training, which improves the model’s ability to understand attributes, relations, and order significantly, leading to better performance on tasks that require compositional understanding. Unlike these approaches, our objective is to investigate the inherent skewed perception embedded in all text embeddings (including target class text and spurious attribute text) and explore the possibilities of using visual representations instead to distill the task-specific core features from VLMs like CLIP for downstream tasks, without the need for group annotations.
@@ 3 Preliminaries
In this study, we explore the spurious features inherent in the CLIP visual representations and assess their impact on classification performance. For a given classification task, we have samples , where represents the input features, as the class labels, as the spurious attributes, and as the group labels. We examine scenarios of distribution shifts occurring between samples across different groups but from same the same class. In the Waterbirds dataset [18], we define {landbird, waterbird}, {land background, water background}, and {landbird on land (), landbird on water (), waterbird on land (), waterbird on water ()}. Notably, and are the minority groups (fewer training samples), whereas and are the majority groups. In the default CelebA dataset [13], the categories are {non-blond hair, blond hair}, {female, male}, and {non-blond hair female (), blond hair female (), non-blond hair male (), blond hair male ()}. With regard to CelebA, are the majority groups, with being the minority group.
@@
4 A Linear Probe Can Achieve Optimal Task Performance
In modern machine learning system design, the goal is often to enhance foundational models with specialized modules for specific downstream tasks. For classification tasks in particular, the ideal is to utilize the same representations derived from a pre-trained model across various classification challenges. Given the nature of spurious correlations — where a feature deemed spurious for one task may be essential for another — we expect the VLMs to capture a broad spectrum of nuanced visual information, and removing the spurious feature by specialized modules. This section delves into the presence of spurious correlations within VLMs and explores whether a simple linear probe can deliver optimal performance.
4.1 Unraveling Spurious Correlations in VLMs
In this section, we investigate the existence of spurious features within CLIP through a comparative analysis of zero-shot binary classification performance on the Waterbirds dataset. We use the text encoder in CLIP to obtain representations for the text labels (“a photo of a landbird” and “a photo of a waterbird”) and the image encoder to extract features from the images. The most similar label, which the model would assign to the image later, is determined by calculating the cosine similarity between the image and text features. To show that spurious correlation impairs the performance of CLIPs, we conduct two experiments: zero-shot classification using the original Waterbirds dataset (with natural background) and using a modified version of the Waterbirds dataset from which the background have been erased based on mask (available in Waterbird dataset). Figure 1 shows one example of the two version of Waterbird dataset.
Figure 2 presents the group accuracy changes across these two scenarios. In the first scenario (using original data), the models exhibit uneven accuracies across majority and minority groups, reflecting the unbalanced group robustness across the dataset. When the backgrounds are removed, the accuracy in recognizing and (minority groups) boosts, as the confounding elements are no longer able to mislead the model. Despite the relatively small change on majority , we see consistent accuracy drop in majority , across three different architectures. This indicates that the model depends more heavily on background information when classifying waterbird-on-water samples compared to landbird-on-land samples. Conclusion 1. The disparity in performance and drastic accuracy change in both minority and majority group, confirms our hypothesis that visual representations in current models are entangled with spurious features that significantly impair classification performance. This raises an fundamental question: does this imply that we are unable to achieve flawless task execution on CLIP representations when faced with spurious features? If not, how?
@@ 4.2 Assessing the Expressiveness of CLIP’s Visual Representations under Linear Probing
In order to see the upper limit of CLIP visual representation with linear transformation, we applied DFR [11], a state-of-the-art method that retrains the last linear layer of the image classifiers (e.g. ResNet) by upweighting the underrepresented samples in the loss function, to 29 attribute classification challenges on the CelebA dataset, where each attribute demonstrated a gender-biased distribution. The size of each group w.r.t. each attributes can be found in supplementary material. Likewise, they also involve four groups based on gender and attribute presence: female without [attribute] (), male without [attribute] (), female with [attribute] (), and male with [attribute] (). Following DFR’s implementation, a linear layer is attached to the CLIP image encoder to facilitate binary classification, with updates restricted solely to the weights of the linear layer. As a supervised method (knowing the group label of each sample), DFR usually signifies the peak performance that a linear layer can attain and reduce the impact of spurious attributes without altering the primary network. High accuracy in these groups suggests that the standard CLIP image encoder successfully captures essential task-related features, not merely relying on these features for predictions under ERM.
Figure 3 showcases the outcomes for various binary attribute classifications with a CLIP ViT-L/14 image encoder and a linear projection head, sorted by ascending average group accuracy (a weighted average determined by the size of each group). The spectrum of attributes ranged from subtle features like straight hair and narrow eyes to more overt characteristics such as eyeglasses, baldness, and hats. Notably, the DFR approach strategy enabled a majority of the attributes — over 25 out of 29 — to achieve more than 75% worst group accuracy (WGA), with more than 23 attribute classification tasks surpassing 80% average group accuracy. The five attributes with the highest accuracies exceeded 95% in both the worst and average group accuracy measures. These results underscore how fine-grained the CLIP’s visual representations are, capturing a comprehensive spectrum of visual information, including subtle features that are typically challenging for human perception. Conclusion 2. Visual representations learned by CLIP are adept at extracting nuanced features within images for various tasks by linear transformation.
5 Language Representations are More Noisy than Expected
VLMs like CLIP prevail partly because of their capability to perform zero-shot inferences guided by intuitive language cues. As demonstrated in the previous section, techniques such as DFR guide us toward identifying an optimal linear probe that can effectively discern variations in core features. Similarly, in zero-shot learning, the linear probe is formed by concatenating text representations. This section explores the biases inherent in zero-shot classification prompts.
@@ 5 Language Representations are More Noisy than Expected
To test this hypothesis, we examined various prompts and corresponding image pairs, calculating the cosine similarity between them. In Figure 4, we evaluated pairs like (“a photo of a camel”/“a photo of a cow”, desert/pasture images), (“a photo of a polar bear”/“a photo of a panda”, glacier/bamboo forest images), and conversely (“a photo of pasture”/“a photo of desert”, camel/cow images), (“a photo of bamboo forest”/“a photo of glacier”, polar bear/panda images), with ensuring the images tested here did not contain the objects mentioned in the prompts. This methodology helps quantify the extent of spurious features embedded in text representations.
@@
Figure 5 shows the result. Notably, the cosine similarity distributions, indicated by the mean (green triangle) and median (orange line), reveal strong correlations—for example, the prompt “a photo of a camel” with camel-free desert images (top left in Figure 4) and the prompt “a photo of a cow” with cow-free pasture images (top left in Figure 4). This pattern is consistent across various tested pairs, underscoring the substantial presence of context-related features in CLIP text representations that are not explicitly present in the prompts. For the prompt and image pair with less pronounced correlations, like “a photo of a dog” to forest and desert, we did not observe the same level of disparity in mean and median value, see supplementary materials for more details.
@@
Conclusion 3. Using text representations for zero-shot classification or debiasing with the representations from spurious attribute prompts [3] could lead to unexpected outcomes, due to the embedded non-target features.
6 Visual Representations Can Refine the Skewed Perception in VLMs
Due to the noisy nature of the text embeddings from CLIP, we wonder if we can use images to construct an optimal linear probe. The availability of background images in the Waterbirds dataset inspires us to explore whether we can use the background images to remove non-core features from the representations of the original Waterbirds images. In Figure 6, we demonstrate our proposed framework VisualDistiller, which achieves high group accuracy with only ERM and a simple projection. We opted for an ERM linear probe (trained by only one epoch, not knowing the group labels of training samples) over zero-shot text classification due to the existence of context-related features in text representations, which compromise classification reliability.
The process is as follows: After encoding the target image via CLIP, we obtained the target image feature . Prior to projection, we aim to isolate the “background” component from . We model this as a linear problem by constructing a subspace in , spanned by “background” vectors . Assuming , where is closest vector to and lies in the orthogonal complement , we define as an matrix of linearly independent columns () and . The orthogonal component is calculated as:
(see proof in supplementary materials). is then processed through the ERM-trained linear probe to produce the final classification result. For simplicity and ease of implementation, we assume Euclidean orthogonality, as using non-Euclidean orthogonality—specifically, employing the inverse of the covariance matrix in the neighborhood space as the metric for inner product calculations—did not yield significant improvement.
The definition of “background” image varies with the dataset. For the Waterbirds dataset, artificially created with images from Places [29] and CUB [23], we defined a range of “background” conditions from least related (random images from Places) to most related (natural environments like lakes and forests) to specific backgrounds used in Waterbirds. Figure 7 shows examples of background images with varying levels of semantic similarity to the Waterbird sample image’s background, which is used in Table 1. In Table 1, we demonstrate that the experiments using semantically more similar “background” images yields higher WGA in Waterbirds. Besides, transitioning to an ERM-trained linear probe enhances WGA further by focusing more sharply on core features, unlike the “contaminated” text representations from CLIP. Both supervised (knowing the corresponding “background” category, denoted by in Table 1) and unsupervised (without knowing the corresponding “background” category, denoted by ) projections were explored, with the supervised setup serving to illustrate the upper limit of this approach, rather than its practical applicability. Increasing the number of “background” vectors generally improves WGA, but with diminishing returns. The VisualDistiller can achieve the WGA of 82.40% without knowing the background image category (20 random images from nature, ViT), only a few points from supervised DFR’s 85.67% performance.